please help me solve the system of equations
x+9y=2 (1)
x=9-9y (2)
answer is ordered pair or there is no solution

Answer by Alan3354(69443)

x+9y=2 (1)
x=9-9y (2) --> x+9y = 9
These are inconsistent, it can't equal 2 and 9. No solution.